If a wire of uniform cross section is stretched such that ,its length is increased by 2%,50%then the %change in its resistance is___
1)4%,100%
2)1%,25%
3)4%,125%
4)4%,50%

Grade:12

1 Answers

Vikas TU
7 years ago
As nothing has been mentioned about area x length = constant we will follow the conventionas:
Resistance is directly proprtional to square of length of Resistance wire.
Thus 
R = kl^2.........................(1)
After stretching,
l becomes =  >0.02l + l = 1.02l
Therfore
R’ = k(1.02l)^2.....................(2)
Divide the eqns. (1) and (2)
R/R’ =  1/(1.02)^2
 
Simlarly do it for 50%.